# Attributing Response to Context: A Jensen–Shannon Divergence Driven Mechanistic Study of Context Attribution in Retrieval-Augmented Generation

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) leverages large language models (LLMs) combined with external contexts to enhance the accuracy and reliability of generated responses. However, reliably attributing generated content to specific context segments, context attribution, remains challenging due to the computationally intensive nature of current methods, which often require extensive fine-tuning or human annotation. In this work, we introduce a novel Jensen–Shannon Divergence driven method to Attribute Response to Context (ARC-JSD), enabling efficient and accurate identification of essential context sentences without additional fine-tuning or surrogate modelling. Evaluations on a wide range of RAG benchmarks, such as TyDi QA, Hotpot QA, and Musique, using instruction-tuned LLMs in different scales demonstrate superior accuracy and significant computational efficiency improvements compared to the previous surrogate-based method. Furthermore, our mechanistic analysis reveals specific attention heads and multilayer perceptron (MLP) layers responsible for context attribution, providing valuable insights into the internal workings of RAG models. Our code is available at https://github.com/ruizheliUOA/ARC_JSD.
@@ 1 Introduction
Nevertheless, verifying the extent to which generated responses are genuinely grounded in their cited context remains a challenging task. Current approaches frequently rely heavily on human annotation (menick2022teaching, ; slobodkin2024attribute, ) or computationally expensive methods such as model fine-tuning and gradient-based feature attribution for accurate attribution (yue2023automatic, ; qi2024model, ; chuang2025selfcite, ), particularly when dealing with extensive documents. For instance, Qi et al. qi2024model utilised distribution shifts between responses generated with and without context to identify relevant tokens and employed gradient-based feature attribution to pinpoint context relevance. Similarly, Chuang et al. chuang2025selfcite enhanced context attribution accuracy through reward-driven fine-tuning within a Direct Preference Optimisation (DPO) framework, based on probability drop and hold analysis of model outputs to context ablation.
@@
To circumvent these computationally intensive methods, Cohen-Wang et al. cohen2024contextcite introduced an inference-time attribution mechanism premised on the assumption that if removing grounded context segments substantially reduces the probability of a generated response, those segments are deemed necessary. Conversely, if retaining only grounded segments maintains response probability, these segments are considered sufficient. By capturing hundreds of probability ablation variations per context-response pair, Cohen-Wang et al. cohen2024contextcite trained a linear surrogate model based on those hundreds of vectors, including the context segment masks and the corresponding generation probability of the original response, to identify context segments crucial for grounding model responses.
@@
However, Cohen-Wang et al cohen2024contextcite still need hundreds of RAG model’s forward calls to collect probability ablation samples for the linear surrogate model training. We propose a novel inference-time Jensen–Shannon Divergence driven method to Attribute Response to Context (ARC-JSD), building upon the inference-attribution assumption above. Our method evaluates the divergence in response distributions generated under the full context compared to sentence-ablated contexts, ranking context sentences based on their JSD differences. This approach offers a significant computational advantage, as it eliminates the need for any additional fine-tuning or surrogate modelling. Furthermore, our ARC-JSD can avoid missing or smoothing non-linearities using JSD to directly quantify actual output distribution shift compared to the linear surrogate modelling cohen2024contextcite .
@@
We empirically evaluate our JSD-driven context attribution approach across multiple question-answering benchmarks, i.e., TyDi QA (clark2020tydi, ), Hotpot QA (yang2018hotpotqa, ), and MuSiQue (trivedi2022musique, ), using state-of-the-art instruction-tuned LLMs including Qwen2-1.5B-Instruct, Qwen2-7B-Instruct (yang2024qwen2, ), Gemma2-2B-Instruct, and Gemma2-9B-Instruct (team2024gemma, ). Our results not only demonstrate improved average accuracy over 10% in context attribution but also achieve computational efficiency, achieving up to a three-fold speedup compared to Cohen-Wang et al. cohen2024contextcite ’s linear-surrogate-based method.
@@
Moreover, we investigate deeper into a mechanistic exploration of context attribution within RAG LLMs by integrating JSD-based analysis with Logit Lens (Nostalgebraist2020, ). Through systematic probing, we identify specific attention heads and multilayer perceptron (MLP) layers critical for context attribution. By subsequently analysing these attention heads and visualising how relevant knowledge is stored in the corresponding MLP layers, we provide concrete evidence of their essential roles in context attribution and further elucidate how contextually relevant information is encoded and utilised within the internal mechanisms of RAG models.
@@
    Developing a lightweight, JSD-driven context attribution method that accurately identifies context sentences critical for grounding generated responses without requiring fine-tuning or surrogate modelling.
@@
    Proposing a versatile, computationally efficient solution that can be readily integrated into any existing RAG-based LLM frameworks and improve RAG model trustworthiness.
@@
    Conducting a detailed mechanistic analysis of RAG LLMs, systematically uncovering and validating specific attention heads and MLP layers responsible for context attribution behaviours.
@@ 2 Related Work
Prior works for context attribution mainly focus on teaching RAG LLMs to generate self-citations for responses, such as few-shot in-context learning (gao2023enabling, ), instruction fine-tuning (ye2024effective, ). Some post-hoc works (chen2023purr, ; qi2024model, ) used an auxiliary language model or gradient-based feature attribution to locate relevant context segments. In general, those methods for context attribution are corroborative (worledge2024unifying, ) in nature, as citations within context are evaluated on whether they support or imply a generated response. Meanwhile, Cohen-Wang et al.; Chuang et al. cohen2024contextcite ; chuang2025selfcite including our work focus on the contributive attribution methods, which are used to identify whether citations cause RAG LLMs to generate a response. Chuang et al. chuang2025selfcite proposed a reward-based fine-tuning with DPO to guide RAG LLMs for context attribution, and Cohen-Wang et al. cohen2024contextcite further trained a linear surrogate model to identify context segments crucial for grounding model responses. However, compared to cohen2024contextcite ; chuang2025selfcite and corroborative methods above, our ARC-JSD method eliminates the need for any additional fine-tuning or surrogate modelling, and it can be directly integrated into any existing RAG-based LLMs.
@@
Existing mechanistic studies mainly focus on the next token generation task to analyse the internal mechanisms of attention heads or MLPs, such as hallucination detection (ferrando2025do, ), multiple-choice questions (li2024anchored, ; wiegreffe2025answer, ; wang2025eliminating, ) and knowledge editing (meng2022locating, ; meng2023massediting, ; katz2024backward, ). Recently, Sun et al. sun2025redeep used a mechanistic interpretability method to analyse attention heads and MLPs of RAG LLMs for the hallucination detection task. Compared to sun2025redeep focusing on locating sources which leads to hallucinations, our proposed ARC-JSD can be regarded as a complementary method to locate citations within context segments and analyse attentions and MLPs, which causes RAG LLMs to generate a correct response. Wu et al. wu2025retrieval focuses on mechanistically analysing retrieval attention heads of RAG LLMs under the Needle-in-the-Haystack (NIAH) setting, where they mainly evaluate whether retrieval attention heads conduct a copy-and-paste operation for retrieving a semantically irrelevant “needle” sentence from the context to the model’s outputs. Compared to wu2025retrieval , which restricts their mechanistic analysis to the NIAH setting where the model performs copy-and-paste retrieval, our work investigates how RAG LLMs mechanistically generate responses based on retrieved content through paraphrasing and contextual integration. This setting better reflects real-world RAG applications [^2] Compared to the traditional RAG to directly map context and response based on their word embeddings, our work has a more general setting, which avoids potential embedding mismatch due to the common paraphrase of RAG LLMs., where models rarely copy text exactly but instead synthesise and rephrase information from retrieved sources.
@@ 3 Background
Consider an autoregressive Transformer-based language model (LLM), denoted as $\mathcal{P}_{\text{LM}}(\cdot)$. Under RAG settings, this model generates responses ($\mathcal{R}$) based on an input query ($\mathcal{Q}$) and associated context ($\mathcal{C}$). Formally, the response generation process can be described as $\mathcal{R}\sim\mathcal{P}_{\text{LM}}(\cdot|\mathcal{C},\mathcal{Q})$, where the context $\mathcal{C}$ consists of sentences $(c_{1},c_{2},\dots,c_{|\mathcal{C}|})$, the query $\mathcal{Q}$ comprises tokens $(q_{1},q_{2},\dots,q_{|\mathcal{Q}|})$, and the generated response $\mathcal{R}$ includes tokens $(r_{1},r_{2},\dots,r_{|\mathcal{R}|})$. Our analysis of context attribution focuses on how the entire response distribution changes when conditioned on the full context set and ablated context alongside the query:
$\mathcal{R}\sim\mathcal{P}_{\text{LM}}(\cdot|c_{1},\dots,c_{|\mathcal{C}|},\mathcal{Q}),\>\mathcal{R}\sim\mathcal{P}_{\text{LM}}(\cdot|\mathcal{C}_{\text{ABLATE}}(c_{i}),\mathcal{Q})\,\text{where}\,\mathcal{C}_{\text{ABLATE}}(c_{i})=\mathcal{C}\setminus\{c_{i}\},\>i\in\{1,\dots,|\mathcal{C}|\}$
Internal Mechanisms of LLMs.
Given a context-query pair represented by a sequence of $T$ tokens $(t_{1},\dots,t_{T})$ drawn from a vocabulary $\mathcal{V}$, tokens are initially encoded into $d$-dimensional vectors $\mathbf{x}_{i}^{0}\in\mathbb{R}^{d}$ through an embedding matrix $W_{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$.
An LLM typically consists of $L$ layers, each composed of attention and MLP modules. These modules iteratively transform token embeddings into residual streams at each layer, denoted $(\mathbf{x}_{1}^{\ell},\dots,\mathbf{x}_{T}^{\ell})$, where $\mathbf{x}_{i}^{\ell}\in\mathbb{R}^{d}$ represents the embedding of token $i$ at layer $\ell$. The residual streams serve as central hubs where attention and MLP modules read and write representations (elhage2021mathematical, ), following the update rule: $\mathbf{x}_{i}^{\ell}=\mathbf{x}_{i}^{\ell-1}+\mathbf{a}_{i}^{\ell}+\mathbf{m}_{i}^{\ell}$, where $\mathbf{a}_{i}^{\ell}$ and $\mathbf{m}_{i}^{\ell}$ denote contributions from the attention and MLP modules at layer $\ell$, respectively. At the final layer $L$, the next token prediction probability distribution is computed as: $\mathcal{P}_{\text{LM}}(t_{T+1}|t_{1:T})=\text{Softmax}\left(W_{U}\sigma(\mathbf{x}_{T}^{L})\right)$, where $W_{U}\in\mathbb{R}^{d\times|\mathcal{V}|}$ is the unembedding matrix, and $\sigma(\cdot)$ denotes pre-unembedding layer normalization.
The attention module, including multiple attention heads, primarily modifies each token’s residual stream representation by attending to prior tokens simultaneously: $\mathbf{a}_{i}^{\ell}=\sum_{h=0}^{H}\mathbf{a}_{i}^{\ell,h}$, where $\mathbf{a}_{i}^{\ell,h}$ indicate each attention head contribution to the residual stream at layer $\ell$. Further details of attention mechanisms are provided in Appendix C.
The MLP modules are often conceptualised as key-value memory structures (geva2021transformer, ; elhage2021mathematical, ; geva2022transformer, ; dai2022knowledge, ). In these modules, columns of $W_{\text{in}}^{\ell}$ serve as keys, while rows of $W_{\text{out}}^{\ell}$ act as corresponding values. The input $\mathbf{x}_{i}^{\ell-1}$ generates a coefficient vector $\mathbf{k}_{i}^{\ell}=\gamma(W_{\text{in}}^{\ell}\mathbf{x}_{i}^{\ell-1})\in\mathbb{R}^{d_{\text{m}}}$ to weight the associated values in $W_{\text{out}}^{\ell}$: $\mathbf{m}_{i}^{\ell}=\sum_{n=1}^{d_{\text{m}}}\mathbf{k}_{i}^{\ell,n}\mathbf{v}_{i}^{\ell,n}$. Further explanations of the MLP mechanisms are available in Appendix C.
@@
Logit lens (Nostalgebraist2020, ) is a mechanistic interpretability method designed to analyse intermediate representations within autoregressive Transformers. Given the LLM architecture described above, logit lens leverages intermediate representations to quantify the direct contribution of attention heads ($\mathbf{a}_{i}^{\ell,h}$), MLP outputs ($\mathbf{m}_{i}^{\ell}$), and residual streams ($\mathbf{x}_{i}^{\ell}$) to token logits: $\text{logit}_{i}^{\ell,h}(\mathbf{a}_{i}^{\ell,h})=W_{U}\sigma(\mathbf{a}_{i}^{\ell,h}),\text{logit}_{i}^{\ell}(\mathbf{m}_{i}^{\ell})=W_{U}\sigma(\mathbf{m}_{i}^{\ell}),\text{logit}_{i}^{\ell}(\mathbf{x}_{i}^{\ell})=W_{U}\sigma(\mathbf{x}_{i}^{\ell})$. Thus, logit lens serves as a powerful tool for pinpointing specific model components crucial to prediction behaviours.
Figure 1: This framework demonstrates how our ARC-JSD works: (a) a RAG LLM $\mathcal{P}_{\text{LM}}(\cdot)$ first generates response $\mathcal{R}$ conditioned on full context $\mathcal{C}$ and query $\mathcal{Q}$ input; (b) By ablating single context sentence once a time, we can calculate probability distribution of the same response $\mathcal{R}$ conditioned on the ablated context $\mathcal{C}_{\text{ABLATE}}(c_{i})$ and query $\mathcal{Q}$; (c) We further calculate the JSD scores about probability distribution of the same response $\mathcal{R}$ conditioned on the full context and ablated context, and locate the most relevant context sentence supporting $\mathcal{R}$ with the highest JSD score.

### v4

# Workshop on Mechanistic Interpretability

@@ Abstract
Retrieval-Augmented Generation (RAG) leverages large language models (LLMs) combined with external contexts to enhance accuracy and reliability of generated responses. However, reliably attributing generated content to specific context segments, context attribution, remains challenging due to computationally intensive nature of current methods, which often require extensive fine-tuning or human annotation. In this work, we introduce a novel Jensen–Shannon Divergence driven method to Attribute Response to Context (ARC-JSD), enabling efficient and accurate identification of essential context sentences without additional fine-tuning, gradient-calculation or surrogate modelling. Evaluations on a wide range of RAG benchmarks, such as TyDi QA, Hotpot QA, and Musique, using instruction-tuned LLMs in different scales demonstrate superior accuracy and significant computational efficiency improvements compared to the previous baselines. Furthermore, our mechanistic analysis reveals specific attention heads and multilayer perceptron (MLP) layers responsible for context attribution, providing valuable insights into the internal workings of RAG models and how they affect RAG behaviours. Code is available at https://github.com/ruizheliUOA/ARC_JSD.
@@ 1 Introduction
Nevertheless, verifying the extent to which generated responses are genuinely grounded in their cited context remains a challenging task. Current approaches frequently rely heavily on human annotation (menick2022teaching, ; slobodkin2024attribute, ) or computationally expensive methods such as model fine-tuning and gradient-based feature attribution for accurate attribution (yue2023automatic, ; qi2024model, ; chuang2025selfcite, ), particularly when dealing with extensive documents. For instance, qi2024model utilised distribution shifts between responses generated with and without context to identify relevant tokens and employed gradient-based feature attribution to pinpoint context relevance. Similarly, chuang2025selfcite enhanced context attribution accuracy through reward-driven fine-tuning within a Direct Preference Optimisation (DPO) framework, based on probability drop and hold analysis of model outputs to context ablation.
@@
To circumvent these computationally intensive methods, cohen2024contextcite introduced an inference-time attribution mechanism premised on the assumption that if removing grounded context segments substantially reduces the probability of a generated response, those segments are deemed necessary. Conversely, if retaining only grounded segments maintains response probability, these segments are considered sufficient. By capturing hundreds of probability ablation variations per context-response pair, cohen2024contextcite trained a linear surrogate model based on those hundreds of vectors, including the context segment masks and the corresponding generation probability of the original response, to identify context segments crucial for grounding model responses.
@@
However, cohen2024contextcite still need hundreds of RAG model’s forward calls to collect probability ablation samples for the linear surrogate model training. We propose a novel inference-time Jensen–Shannon Divergence driven method to Attribute Response to Context (ARC-JSD), building upon the inference-attribution assumption above. Our method evaluates the divergence in response distributions generated under the full context compared to sentence-ablated contexts, ranking context sentences based on their JSD differences because of JSD’s symmetric, finite, scale-free, and bounded properties (see § 3 and § 8 for details). This approach offers a significant computational advantage, as it eliminates the need for any additional fine-tuning or surrogate modelling. Furthermore, our ARC-JSD can avoid missing or smoothing non-linearities using JSD to directly quantify actual output distribution shift compared to the linear surrogate modelling (cohen2024contextcite, ).
@@
We empirically evaluate our JSD-driven context attribution approach across multiple question-answering benchmarks, i.e., TyDi QA (clark2020tydi, ), Hotpot QA (yang2018hotpotqa, ), and MuSiQue (trivedi2022musique, ), using state-of-the-art instruction-tuned LLMs including Qwen2-1.5B-Instruct, Qwen2-7B-Instruct (yang2024qwen2, ), Gemma2-2B-Instruct, and Gemma2-9B-Instruct (team2024gemma, ). Our results not only demonstrate improved average accuracy over 10% in context attribution but also achieve computational efficiency, achieving up to a three-fold speedup compared to cohen2024contextcite ’s linear-surrogate-based and other gradient-based baselines.
@@
Moreover, we investigate deeper into a mechanistic exploration of context attribution within RAGs by integrating JSD-based analysis with Logit Lens (Nostalgebraist2020, ). Through systematic probing, we identify specific attention heads and MLP layers critical for context attribution. By subsequently manipulating these located attention heads and MLPs using JSD scores as confidence gates, we can control RAGs and further mitigate the hallucination rate. In addition, we can visualise how relevant knowledge is stored in the corresponding MLP layers.
@@
    Developing a JSD-driven context attribution method that accurately identifies context critical for grounding responses without requiring fine-tuning, surrogate modelling or gradient calculation.
@@
    Proposing a computationally efficient solution that can be readily integrated into any existing RAGs. Conducting a detailed mechanistic analysis of RAG, systematically uncovering and validating attention heads and MLP layers responsible for context attribution behaviours.
@@
    Mitigating the hallucination rate and controlling RAG behaviour by manipulating located attention heads and MLPs within RAG.
@@ 2 Related Work
Prior works for context attribution mainly focus on teaching RAG LLMs to generate self-citations for responses, such as few-shot in-context learning (gao2023enabling, ), instruction fine-tuning (ye2024effective, ). Some post-hoc works (chen2023purr, ; qi2024model, ) used an auxiliary language model or gradient-based feature attribution to locate relevant context segments. In general, those methods for context attribution are corroborative (worledge2024unifying, ) in nature, as citations within context are evaluated on whether they support or imply a generated response. Meanwhile, cohen2024contextcite ; chuang2025selfcite ; cohen2025learning ; liu2024attribot including our work focus on the contributive attribution methods, which are used to identify whether citations cause RAG LLMs to generate a response. chuang2025selfcite proposed a reward-based fine-tuning with DPO to guide RAG LLMs for context attribution, and cohen2024contextcite ; cohen2025learning further trained a linear surrogate model to identify context segments crucial for grounding model responses. liu2024attribot focuses on formalising and comparing different attribution acceleration methods and ignores attribution accuracy improvements. However, compared to cohen2024contextcite ; chuang2025selfcite ; cohen2025learning and corroborative methods above, our ARC-JSD method eliminates the need for any additional fine-tuning or surrogate modelling, and it can be directly integrated into any existing RAGs.
@@
Existing mechanistic studies focus on next token generation task to analyse internal mechanisms of attention heads or MLPs, such as hallucination detection (ferrando2025do, ), multiple-choice questions (li2024anchored, ; wiegreffe2025answer, ; wang2025eliminating, ) and knowledge editing (meng2022locating, ; meng2023massediting, ; katz2024backward, ). Recently, sun2025redeep used a mechanistic interpretability method to analyse attention heads and MLPs of RAGs for the hallucination detection task. Compared to sun2025redeep focusing on locating sources which leads to hallucinations, our proposed ARC-JSD can be regarded as a complementary method to locate citations within context segments and analyse attentions and MLPs, which causes RAG LLMs to generate a correct response. wu2025retrieval focuses on mechanistically analysing retrieval attention heads of RAG LLMs under the Needle-in-the-Haystack (NIAH) setting, where they mainly evaluate whether retrieval attention heads conduct a copy-and-paste operation for retrieving a semantically irrelevant “needle” sentence from the context to the model’s outputs. Compared to wu2025retrieval , which restricts their mechanistic analysis to the NIAH setting where the model performs copy-and-paste retrieval, our work investigates how RAG LLMs mechanistically generate responses based on retrieved content through paraphrasing and contextual integration. This setting better reflects real-world RAG applications [^2] Compared to traditional RAG to directly map prompts based on their word embeddings, our work has a more general setting, which avoids potential embedding mismatch due to the common paraphrase of RAGs., where models rarely copy text exactly but instead synthesise and rephrase information from retrieved sources.
Figure 1: This framework demonstrates how our ARC-JSD works: (a) a RAG LLM $\mathcal{P}_{\text{LM}}(\cdot)$ first generates response $\mathcal{R}$ conditioned on full context $\mathcal{C}$ and query $\mathcal{Q}$ input; (b) By ablating single context sentence once a time, we can calculate probability distribution of the same response $\mathcal{R}$ conditioned on the ablated context $\mathcal{C}_{\text{ABLATE}}(c_{i})$ and query $\mathcal{Q}$; (c) We further calculate JSD scores about probability distribution of the same response $\mathcal{R}$ conditioned on full context and ablated context, and locate the most relevant context sentence supporting $\mathcal{R}$ with the highest JSD score. Then, we apply JSD-based metric to internal components of RAGs: (d) For each attention head or MLP output at each layer, we calculate probability distribution of the same response $\mathcal{R}$ conditioned on the same query $\mathcal{Q}$ with full context $\mathcal{C}$ and ablated context $\mathcal{C}_{\text{ABLATE}}(c_{\text{top-1}})$ by removing top relevant context sentence based on § 4.1; (e) We can further locate top-NN relevant attention heads or MLPs which contribute the context attribution by ranking the collected JSD scores with a descending order.
@@ 3 Background
Consider an autoregressive Transformer-based language model (LLM), denoted as $\mathcal{P}_{\text{LM}}(\cdot)$. Under RAG settings, this model generates responses ($\mathcal{R}$) based on an input query ($\mathcal{Q}$) and associated context ($\mathcal{C}$). Formally, response generation process can be described as $\mathcal{R}\sim\mathcal{P}_{\text{LM}}(\cdot|\mathcal{C},\mathcal{Q})$, where context $\mathcal{C}$ consists of sentences $(c_{1},c_{2},\dots,c_{|\mathcal{C}|})$, the query $\mathcal{Q}$ comprises tokens $(q_{1},q_{2},\dots,q_{|\mathcal{Q}|})$, and the generated response $\mathcal{R}$ includes tokens $(r_{1},r_{2},\dots,r_{|\mathcal{R}|})$. Our analysis of context attribution focuses on how the entire response distribution changes when conditioned on the full context set and ablated context alongside the query: $\mathcal{R}\sim\mathcal{P}_{\text{LM}}(\cdot|c_{1},\dots,c_{|\mathcal{C}|},\mathcal{Q}),\>\mathcal{R}\sim\mathcal{P}_{\text{LM}}(\cdot|\mathcal{C}_{\text{ABLATE}}(c_{i}),\mathcal{Q})\,\text{where}\,\mathcal{C}_{\text{ABLATE}}(c_{i})=\mathcal{C}\setminus\{c_{i}\},\>i\in\{1,\dots,|\mathcal{C}|\}$.
@@
Logit lens (Nostalgebraist2020, ) is a mechanistic interpretability method designed to analyse intermediate representations within autoregressive Transformers. Given the LLM architecture described in Appendix D, logit lens leverages intermediate representations to quantify the direct contribution of attention heads ($\mathbf{a}_{i}^{\ell,h}$), MLP outputs ($\mathbf{m}_{i}^{\ell}$), and residual streams ($\mathbf{x}_{i}^{\ell}$) to token logits: $\text{logit}_{i}^{\ell,h}(\mathbf{a}_{i}^{\ell,h})=W_{U}\sigma(\mathbf{a}_{i}^{\ell,h}),\text{logit}_{i}^{\ell}(\mathbf{m}_{i}^{\ell})=W_{U}\sigma(\mathbf{m}_{i}^{\ell}),\text{logit}_{i}^{\ell}(\mathbf{x}_{i}^{\ell})=W_{U}\sigma(\mathbf{x}_{i}^{\ell})$. Thus, logit lens serves as a powerful tool for pinpointing specific model components crucial to prediction behaviours.
JSD for Context Attribution.
JSD is a symmetrised, smoothed variant of Kullback–Leibler (KL) divergence that quantifies information gap (in bits) between two probability distributions. Because it is symmetric, finite, scale-free, and bounded in $0,\log 2]$, JSD allows scores from different layers to be compared directly, without sensitivity to arbitrary logit shifts. Following “logit-lens” perspective of sun2025redeep , we treat JSD as model’s belief of how much its next-token distribution will change. Concretely, we compute JSD between the full-context token distribution and the distribution obtained after removing a single retrieved sentence $c_{i}$. A high divergence indicates that the model’s internal representations (and therefore its output logits) depend strongly on $c_{i}$. Empirically, ablating the sentence with the highest JSD causes the largest drop in answer likelihood, validating JSD as a concise and reliable signal for context attribution in RAG models (See § 8 for comparisons among JSD, Wasserstein, Total Variation (TV) and Maximum Mean Discrepancy (MMD)).